### v3

@@ 5.1 Discrete Action Space
Suppose $\mathcal{X}_{t}={\mathcal{S}}_{d-1}$ for all $t\in[T]$. Define $\beta^{PVT}_{t}\coloneqq\iota_{t}\beta^{RLS}_{t}$ and
[…]
where we let
$\quad m_{t}=(\|\widehat{\theta}_{t}\|_{V_{t}}^{2}-(\beta^{RLS}_{t})^{2})/(\|\widehat{\theta}_{t}\|+\beta^{RLS}_{t}/\lambda_{d}(V_{t}))^{2},\quad M_{t}=\|\widehat{\theta}_{t}\|_{V_{t}^{2}}^{2}/(\|\widehat{\theta}_{t}\|_{V_{t}}^{2}-(\beta^{PVT}_{t})^{2})$
and $k\in[d]$ is such that $\lambda_{k}(V)\leq M_{t}\leq\lambda_{k+1}(V)$. Then for all $t\in[T]$, conditioned on event $\widehat{\mathcal{A}}_{t}\cap\widetilde{\mathcal{A}}_{t}$, it holds for all $s\leq t$ that
$\inf_{\theta\in\mathcal{E}_{t}^{PVT}}\|x^{\star}_{s}(\theta)\|_{V_{s}^{-1}}\geq\Psi_{s},\quad\sup_{\theta\in\mathcal{E}_{s}^{RLS}}\|x^{\star}_{s}(\theta)\|_{V_{s}^{-1}}\leq\Phi_{s},\quad\alpha_{s}\leq\widehat{\alpha}_{s}\coloneqq\Phi_{s}/\Psi_{s}.$
To better understand what Theorem 2 implies, we discuss some special cases.
Case 1: a pure exploration regime.
When the decision-maker doesn’t care about the regret and adopts a pure exploration algorithm that plays actions in all directions sequentially, we expect $\lambda_{i}(V_{t})=\mathcal{O}(t)$ for all $i=1,\dots,d$. Then $\widehat{\alpha}_{t}\leq C$ for some constant $C>0$. Specifically, if $V_{t}=D\mathbb{I}_{d}$ for some constant $D>0$, we have $\|x\|_{V_{t}^{-1}}=D^{-1}\|x\|=D^{-1}$ for all $x\in{\mathcal{S}}_{d-1}$ and hence $\widehat{\alpha}_{t}=1$
Case 2: linear structures.
When the reward takes a linear form in the action, for example, the stochastic linear bandits we study, it’s observed in practice that the estimate $\widehat{\theta}_{t}$ tends to align in the first eigenspace of the uncertainty structure $V_{t}$ (empirically validated in Appendix 11). To see this, note that in order to maximize the reward in an online manner, bandit algorithms tend to select actions towards the confidence ellipsoid centered at $\widehat{\theta}_{t}$, hence forcing more exploration along the direction of $\widehat{\theta}_{t}$, especially at the late stage of the algorithm when the ellipsoid is small. The following proposition states that, as long as the center $\widehat{\theta}_{t}$ of the confidence ellipsoid $\mathcal{E}_{t}$ tends to be aligned with the first eigenspace of $V_{t}$, the corresponding uncertainty ratio tends to 1, indicating a regret bound matching the minimax $\widetilde{\mathcal{O}}(d\sqrt{T})$ rate by Theorem 1. This provides a plausible explanation for the empirical success of LinTS and Greedy.
Proposition 5
Suppose $\lim_{t\to\infty}\frac{\|\widehat{\theta}_{t}\|_{V_{t}}^{2}}{\|\widehat{\theta}_{t}\|^{2}}=\lambda_{\max}(V_{t})$, it holds that $\lim_{t\to\infty}\widehat{\alpha}_{t}=1$.
@@
Clearly, one can further improve the bound in Equation (6) by plugging in the selected action $\widetilde{x}_{t}$. This is true for standard algorithms such as OFUL, LinTS, and Greedy. Explicitly, the uncertainty ratio can be bounded using $\widetilde{x}_{t}$ and Theorem 2 as
$\alpha_{t}\leq\frac{\sup_{\theta\in\mathcal{E}_{t}^{RLS}}\|x^{\star}_{t}(\theta)\|_{V_{t}^{-1}}}{\|\widetilde{x}_{t}\|_{V_{t}^{-1}}}\leq\frac{\Phi_{t}}{\|\widetilde{x}_{t}\|_{V_{t}^{-1}}}.$ (7)
However, for the variants of POFUL that we are going to propose in the Section 6, it is necessary to calculate the bound before selecting the action $\widetilde{x}_{t}$. Therefore, the bound shouldn’t depend on $\widetilde{x}_{t}$.
5.3 Sketch of the Proof
@@
In this section, we present the proof of Theorem 2. In the remaining part of this section, we use a general confidence ellipsoid $\mathcal{E}_{t}(\delta)\coloneqq\left\{\theta\in\mathbb{R}^{d}:\|\theta-\widehat{\theta}_{t}\|_{V_{t}}\leq\beta_{t}(\delta)\right\}$ to represent $\mathcal{E}^{RLS}_{t}$ and $\mathcal{E}^{PVT}_{t}$, since the proof works for both of them.
@@
6 A Meta-Algorithm for Course-Correction
In this section, we show how the data-driven regret bound can be utilized to improve standard bandit algorithms. Specifically, we propose a meta-algorithm that constructs course-corrected variants of the base algorithms. This ensures they achieve a minimax-optimal frequentist regret guarantee (up to constant factors) while retaining the majority of the original characteristics of the base algorithms, such as computational efficiency and low regret in most cases.
We take LinTS as an example and propose the algorithm Linear Thompson Sampling with Maximum Regret (Proxy) (TS-MR). The idea is to measure the performance of LinTS using $\widehat{\mu}_{t}$ and avoid bad LinTS actions by switching to OFUL actions when necessary.
Specifically, at each time $t$, TS-MR calculates the upper bound $\widehat{\mu}_{t}$ and compares it with a preset threshold $\mu$. If $\widehat{\mu}_{t}>\mu$, LinTS might be problematic and TS-MR takes an OFUL action to ensure a low instantaneous regret; if $\widehat{\mu}_{t}\leq\mu$, TS-MR takes the LinTS action. We remark that setting $\iota_{t}\equiv 0$ yields the corresponding Greedy-MR algorithm. The pseudocode is presented in Algorithm 2.
Algorithm 2 TS-MR (Greedy-MR)
0:  $T$, $\delta$, $\{\iota_{t}\}_{t\in[T]}$, $\mu$
  Initialize $V_{0}\leftarrow\lambda\mathbb{I}$, $\widehat{\theta}_{1}\leftarrow 0$, $\delta^{\prime}\leftarrow\delta/2T$
  for t = 1, 2 , …, T  do
     Calculate $\widehat{\alpha}_{t}$ using Theorem 2
     if $\widehat{\mu}_{t}\leq\mu$ then
        $\widetilde{\theta}_{t}\leftarrow\widehat{\theta}_{t}+\iota_{t}\beta_{t}^{RLS}(\delta^{\prime})V_{t}^{-\frac{1}{2}}\eta_{t}$
     else
        ${x}_{t}\leftarrow\arg\max_{x\in\mathcal{X}_{t}}\arg\max_{\theta\in\mathcal{E}_{t}^{RLS}(\delta^{\prime})}\langle{x},{\widetilde{\theta}_{t}}\rangle$
     end if
     Observe reward $r_{t}$
     $V_{t+1}\leftarrow V_{t}+{x}_{t}{x}_{t}^{\top}$
     $\widehat{\theta}_{t+1}\leftarrow V_{t+1}^{-1}\sum_{s=1}^{t}x_{s}r_{s}$.
  end for
By construction, the course-corrected algorithms have $\mu_{t}\leq\max\{\mu,2\}$ for all $t\in[T]$. Using Theorem 1, we get the following frequentist regret for them, which is optimal up to a constant factor.
Corollary 1
TS-MR and Greedy-MR achieve a frequentist regret of $\mathcal{R}(T)=\widetilde{\mathcal{O}}\left(\max\{\mu,2\}d\sqrt{T}\right)$.
Remark 4
In practice, the choice of the threshold $\mu$ depends on the problem settings. In a high-risk setting where LinTS and Greedy fail with high probability, one can set a small $\mu$ so that TS-MR and Greedy-MR select more OFUL actions to guarantee the necessary amount of exploration. In a low-risk setting where original LinTS and Greedy algorithms work well, one can set a large $\mu$ and hence TS-MR and Greedy-MR select TS and greedy actions respectively to avoid unnecessary exploration and save the computational cost.
7 Simulations
In this section, we present the simulation results of TS-MR, Greedy-MR, and other baseline algorithms.
7.1 Synthetic datasets
(a) example 1
(b) example 2
(c) example 3
Figure 3: Comparison of the cumulative regret incurred by TS-MR and Greedy-MR, versus the baseline algorithms in Examples 1-3.
We fix $d=50$, and sample $\theta^{\star}\sim\text{Unif}(\{\theta\in\mathbb{R}^{d}|\|\theta\|=10\})$ on a sphere with fixed norm. At each time $t$, we generate 100 i.i.d. random actions sampled from $\text{Unif}({\mathcal{S}}_{d-1})$ to form $\mathcal{X}_{t}$. This is a basic example of standard stochastic linear bandit problems without any extra structure. We set the threshold $\mu=8$ for TS-MR and Greedy-MR. We average simulation results over 100 independent runs. The results are summarized in Figure 2(a). TS-Freq shows pessimistic regret due to the inflation of the posterior, while other algorithms in general perform well.
We fix $d=50$, and sample $\theta^{\star}\sim\text{Unif}(\{\theta\in\mathbb{R}^{d}|\|\theta\|=70\})$. At each time $t$, we first generate a random vector $u_{t}\sim\mathcal{N}(0,\mathbb{I}_{5})$ and let $x_{t,i}\in\mathbb{R}^{50}$ be the vector whose $i$-th block of size 5 is a copy of $u_{t}$, and other components are 0. Then $X_{t}=\{x_{t,i}\}_{i\in[10]}$ is an action set of size 10, sharing the same feature $u_{t}$ in different blocks. This problem is equivalent to a 10-armed contextual bandit. We ser $\mu=12$ for TS-MR and Greedy-MR. Again, we average simulation results over 100 independent runs. The results are summarized in Figure 2(b). In this setting, Greedy performs suboptimally due to a lack of exploration for some arms. Nevertheless, Greedy-MR outperforms both Greedy and OFUL by adaptively choosing OFUL actions only when it detects large regret proxy $\widehat{\mu}_{t}$.
Example 3. Prior mean mismatch (Hamidi and Bayati, 2020b).
This is an example in which LinTS is shown to incur linear Bayesian regret. We sample $\theta^{\star}\sim\mathcal{N}(m\mathbf{1}_{3d},\mathbb{I}_{3d})$ and fix the action set $\mathcal{X}_{t}=\{0,x_{a},x_{b}\}$ for all $t\in[T]$, where $x_{a}=-\sum_{i=1}^{d}\frac{e_{i}}{\sqrt{3d}},\quad x_{b}=\sum_{i=11}^{3d}\frac{e_{i}}{\sqrt{3d}}-\sum_{i=1}^{d}\frac{e_{i}}{\sqrt{3d}}$. It is shown in Hamidi and Bayati (2020b) that, when LinTS takes a wrong prior mean as input, it has a large probability to choose $\widetilde{x}_{2}=0$, conditioned on $\widetilde{x}_{1}=x_{a}$. Note that choosing the zero action brings no information update to LinTS, it suffers a linear Bayesian regret when trying to escape from the zero action. We let $m=10$ and set $d=10$, so the problem is a $30$-dimensional linear bandit. We ser $\mu=12$ for TS-MR and Greedy-MR. We carried out 100 independent runs. The results are summarized in Figure 2(c). We see both LinTS and Greedy incur linear regrets as expected, while TS-MR and Greedy-MR, switch to OFUL adaptively to tackle this hard problem and achieve sublinear regret.
7.2 Real-world datasets
We explore the performance of standard POFUL algorithms and the proposed TS-MR and Greedy-MR algorithms on real-world datasets. We consider three datasets for classification tasks on the OpenML platform, Cardiotocography, JapaneseVowels and Segment. They focus on healthcare, pattern recognition, and computer vision problems respectively.
Setup.
We follow the standard approach in the literature that converts classification tasks to contextual bandit problems, and then embeds it into linear bandit problems in the same way as Example 2 in Section 6. Specifically, we regard each class as an action so that at each time, the decision-maker assigns a class to the feature observed and receives a binary reward, namely 1 for assigning the correct class and 0 otherwise, plus a Gaussian noise.
We plot the cumulative regret (averaged over 100 duplicates) for all algorithms. Figure 4 shows that for all real-world datasets: OFUL and TS-Freq perform poorly due to their conservative exploration; LinTS and Greedy are achieving empirical success even though they don’t have theoretical guarantees; TS-MR and Greedy-MR retain the desirable empirical performance of LinTS and Greedy, while enjoying the minimax optimal frequentist regret bound.
(a) Cardiotocography dataset.
(b) JapaneseVowels dataset.
(c) Segment dataset.
Figure 4: Cumulative regret of all algorithms on real-world datasets.
8 Conclusion
In this work, we propose a data-driven framework to analyze the frequentist regret of POFUL, a family of algorithms that includes OFUL, LinTS, TS-Freq, and Greedy as special cases. Our approach allows for the computation of an instance-dependent frequentist regret bound for POFUL during implementation, which subsequently informs the course-correction of the algorithm. Our technique conducts a novel real-time geometric analysis of the $d$-dimensional confidence ellipsoid to fully leverage the historical information and might be of independent interest. As applications, we propose TS-MR and Greedy-MR algorithms that enjoy provable minimax optimal frequentist regret and demonstrate their ability to adaptively switch to OFUL when necessary in hard problems where LinTS and Greedy fail. We hope this work provides a steady step towards bridging the gap between theoretical guarantees and empirical performance of bandit algorithms such as LinTS and Greedy.
{APPENDICES}

### v5

@@ 5.1 Continuous action space
Suppose $\mathcal{X}_{t}={\mathcal{S}}_{d-1}$ for all $t\in[T]$. Define $m_{t}=(\|\widehat{\theta}_{t}\|_{V_{t}}^{2}-(\beta^{RLS}_{t,\delta^{\prime},\lambda_{\text{reg}}})^{2})/(\|\widehat{\theta}_{t}\|+\beta^{RLS}_{t,\delta^{\prime},\lambda_{\text{reg}}}/\lambda_{d}(V_{t}))^{2}$, $M_{t}=\|\widehat{\theta}_{t}\|_{V_{t}^{2}}^{2}/(\|\widehat{\theta}_{t}\|_{V_{t}}^{2}-(\beta^{PVT}_{t,\delta^{\prime},\lambda_{\text{reg}}})^{2})$. Let $k\in[d]$ be the integer that satisfies $\lambda_{k}(V)\leq M_{t}\leq\lambda_{k+1}(V)$. Define $\beta^{PVT}_{t,\delta^{\prime},\lambda_{\text{reg}}}\coloneqq\iota_{t}\beta^{RLS}_{t,\delta^{\prime},\lambda_{\text{reg}}}$ and
$\Phi_{t}$ =$=\begin{cases}(\lambda_{1}^{-1}(V_{t})+\lambda_{d}^{-1}(V_{t})-m_{t}\lambda_{1}^{-1}(V_{t})\lambda_{d}^{-1}(V_{t}))^{\frac{1}{2}},&\ \ \ \ \textit{if}\ \|\widehat{\theta}_{t}\|_{V_{t}}\geq\beta^{RLS}_{t,\delta^{\prime},\lambda_{\text{reg}}}\\ \lambda_{d}^{-\frac{1}{2}}(V_{t}),&\ \ \ \ \textit{if}\ \|\widehat{\theta}_{t}\|_{V_{t}}<\beta^{RLS}_{t,\delta^{\prime},\lambda_{\text{reg}}}\end{cases},$
$\Psi_{t}$ =$=\begin{cases}(\lambda_{k}^{-1}(V_{t})+\lambda_{k+1}^{-1}(V_{t})-M_{t}\lambda_{k}^{-1}(V_{t})\lambda_{k+1}^{-1}(V_{t}))^{\frac{1}{2}},&\textit{if}\ \|\widehat{\theta}_{t}\|_{V_{t}}\geq\beta^{PVT}_{t,\delta^{\prime},\lambda_{\text{reg}}}\\ \lambda_{1}^{-\frac{1}{2}}(V_{t}),&\textit{if}\ \|\widehat{\theta}_{t}\|_{V_{t}}<\beta^{PVT}_{t,\delta^{\prime},\lambda_{\text{reg}}}\end{cases}.$
Then for all $t\in[T]$, conditioned on $\widehat{\mathcal{A}}_{t}\cap\widetilde{\mathcal{A}}_{t}$, it holds for all $s\leq t$ that $\alpha_{s}\leq\widehat{\alpha}_{s}\coloneqq\Phi_{s}/\Psi_{s}$.
To better understand what Theorem 2 implies, we discuss some special cases in Appendix 14 and provide empirical validations for them.
6 A Meta-Algorithm for Course-Correction
This section demonstrates how the data-driven regret bound can enhance standard bandit algorithms. We propose a meta-algorithm that creates course-corrected variants of base algorithms, achieving minimax-optimal frequentist regret guarantees while preserving most original characteristics, including computational efficiency and typically low regret.
We take LinTS as an example of the base algorithm, and propose the algorithm Linear Thompson Sampling with Maximum Regret (Proxy) (TS-MR). The idea is to measure the performance of LinTS using $\widehat{\mu}_{t}$ and avoid bad LinTS actions by switching to OFUL actions. Specifically, at each time $t$, TS-MR calculates the upper bound $\widehat{\mu}_{t}$ and compares it with a preset threshold $\mu$. If $\widehat{\mu}_{t}>\mu$, LinTS might be problematic and TS-MR takes an OFUL action to ensure a low instantaneous regret; if $\widehat{\mu}_{t}\leq\mu$, TS-MR takes the LinTS action. We remark that setting $\iota_{t}=0$ for all $t\in[T]$ yields the corresponding Greedy-MR algorithm. The pseudocode is presented in Algorithm 2 in Appendix 15.
@@ 6 A Meta-Algorithm for Course-Correction
Computing $\widehat{\mu}_{t}$ primarily requires SVD decomposition of the sample covariance matrix $V_{t}$. Since $V_{t}=\lambda_{\text{reg}}I_{d}+\sum_{s=1}^{t}x_{s}x_{s}^{\top}$ is updated via rank-one matrices, its SVD can be efficiently updated (Gandhi and Rajgor, 2017), preventing computational bottlenecks.
By design, course-corrected algorithms ensure that $\mu_{t}\leq\max\{\mu,2\}$ for all $t\in[T]$. Substituting this upper bound into Theorem 1 establishes that these algorithms achieve optimal frequentist regret, up to a constant factor.
Corollary 1
TS-MR and Greedy-MR achieve a frequentist regret of $\widetilde{\mathcal{O}}(\max\{\mu,2\}d\sqrt{T})$.
In high-risk settings where LinTS and Greedy may fail, a small $\mu$ ensures TS-MR and Greedy-MR select more OFUL actions, promoting sufficient exploration. Conversely, in low-risk settings where the original algorithms perform well, a large $\mu$ favors TS and greedy actions, minimizing unnecessary exploration and reducing computational cost. In Appendix 16, we show how $\mu$ impacts the fraction of OFUL actions in TS-MR and Greedy-MR and their performance. Results indicate that course-corrected algorithms maintain robustness across a range of moderate $\mu$ values, suggesting that the precise selection of $\mu$ is unlikely to present a significant practical concern.
7 Simulations
We aim to compare TS-MR, Greedy-MR, and key baseline algorithms, via simulation.
7.1 Synthetic datasets
We conduct simulations on three representative synthetic examples. We average simulation results over 100 independent runs for each of the examples. The results are shown in Figure 3.
We fix $d=50$, and sample $\theta^{\star}\sim\text{Unif}(\{\theta\in\mathbb{R}^{d}|\|\theta\|=10\})$ on a sphere with fixed norm. At each time $t$, we generate 100 i.i.d. random actions sampled from $\text{Unif}({\mathcal{S}}_{d-1})$ to form $\mathcal{X}_{t}$. This is a basic example of standard stochastic linear bandit problems without any extra structure. We set the threshold $\mu=8$ for TS-MR and Greedy-MR. TS-Freq shows pessimistic regret due to the inflation of the posterior, while other algorithms in general perform well.
We fix $d=50$, and sample $\theta^{\star}\sim\text{Unif}(\{\theta\in\mathbb{R}^{d}|\|\theta\|=70\})$. At each time $t$, we first generate a random vector $u_{t}\sim\mathcal{N}(0,I_{5})$ and let $x_{t,i}\in\mathbb{R}^{50}$ be the vector whose $i$-th block of size 5 is a copy of $u_{t}$, and other components are 0. Then $X_{t}=\{x_{t,i}\}_{i\in[10]}$ is an action set of size 10, sharing the same feature $u_{t}$ in different blocks. This problem is equivalent to a 10-armed contextual bandit. We set $\mu=12$ for TS-MR and Greedy-MR. In this setting, Greedy performs suboptimally due to a lack of exploration for some arms. Nevertheless, Greedy-MR outperforms both Greedy and OFUL by adaptively choosing OFUL actions only when it detects large regret proxy $\widehat{\mu}_{t}$.
Example 3. Prior mean mismatch (Hamidi and Bayati, 2020b).
This is an example in which LinTS is shown to incur linear Bayesian regret. We sample $\theta^{\star}\sim\mathcal{N}(m\mathbf{1}_{3d},I_{3d})$ and fix the action set $\mathcal{X}_{t}=\{0,x_{a},x_{b}\}$ for all $t\in[T]$, where $x_{a}=-\sum_{i=1}^{d}e_{i}/\sqrt{3d},\quad x_{b}=\sum_{i=11}^{3d}e_{i}\sqrt{3d}-\sum_{i=1}^{d}e_{i}/\sqrt{3d}$. It is shown in Hamidi and Bayati (2020b) that, when LinTS takes a wrong prior mean as input, it has a large probability to choose $\widetilde{x}_{2}=0$, conditioned on $\widetilde{x}_{1}=x_{a}$. Note that choosing the zero action brings no information update to LinTS, it suffers a linear Bayesian regret when trying to escape from the zero action. We let $m=10$ and set $d=10$, so the problem is a $30$-dimensional linear bandit. We set $\mu=12$ for TS-MR and Greedy-MR. We see both LinTS and Greedy incur linear regrets as expected, while TS-MR and Greedy-MR, switch to OFUL adaptively to tackle this hard problem and achieve sublinear regret.
(a) example 1
(b) example 2
(c) example 3
(d) example 1
(e) example 2
(f) example 3
Figure 3: Simulation results on synthetic data. (a) - (c): Cumulative regret of TS-MR and Greedy-MR versus baseline algorithms. Shaded regions show ±$\pm 2$ SE of mean regret. (d) - (f): Fraction of OFUL actions in TS-MR and Greedy-MR.
7.2 Real-world datasets
We explore the performance of standard POFUL algorithms and the proposed TS-MR and Greedy-MR algorithms on real-world datasets. We use three classification datasets from OpenML: Cardiotocography, JapaneseVowels, and Segment, representing healthcare, pattern recognition, and computer vision domains. Following Bietti et al. (2021); Bastani et al. (2021), we convert these classification tasks to contextual bandit problems and embed them into linear bandit problems as in Example 2, Section 7.1. Each class becomes an action where the decision-maker receives a binary reward (1 for correct classification, 0 otherwise) plus Gaussian noise.
We plot the cumulative regret (averaged over 100 runs) for all algorithms. Figure 4 shows that for all real-world datasets: OFUL and TS-Freq perform poorly due to their conservative exploration; LinTS and Greedy are achieving empirical success even though they don’t have theoretical guarantees; TS-MR and Greedy-MR retain the desirable empirical performance of LinTS and Greedy, while enjoying the minimax optimal frequentist regret bound.
(a) Cardiotocography dataset
(b) JapaneseVowels dataset
(c) Segment dataset
(d) Cardiotocography dataset
(e) JapaneseVowels dataset
(f) Segment dataset
Figure 4: Simulation results on real-world datasets. (a) - (c): Cumulative regret of all algorithms. Shaded regions show the ±$\pm 2$ SE of the mean regret. (d) - (f): Fraction of OFUL actions of TS-MR and Greedy-MR.
Remark 4
Simulation results in Figures 3 and 4 show OFUL actions are primarily used in the early stages. This indicates: (1) Greedy-MR and TS-MR implement OFUL actions only when necessary, maintaining a low OFUL fraction throughout most of the time horizon, substantially reducing computational cost; and (2) limited course-corrected exploration at the beginning efficiently remedies TS and Greedy in problematic instances.
8 Conclusion
In this work, we propose a data-driven framework to analyze the frequentist regret of POFUL, a family of algorithms that includes OFUL, LinTS, TS-Freq, and Greedy as special cases. Our approach allows for the computation of a data-driven frequentist regret bound for POFUL during implementation, which subsequently informs the course-correction of the algorithm. Our technique conducts a novel real-time geometric analysis of the $d$-dimensional confidence ellipsoid to fully leverage the historical information and might be of independent interest. As applications, we propose TS-MR and Greedy-MR algorithms that enjoy provable minimax optimal frequentist regret and demonstrate their ability to adaptively switch to OFUL when necessary in hard problems where LinTS and Greedy fail. We hope this work provides a steady step towards bridging the gap between theoretical guarantees and empirical performance of bandit algorithms such as LinTS and Greedy.
{APPENDICES}
9 Sketch of the Proof
@@
In this section, we present the proof of Theorem 2. In the remaining part of this section, we use a general confidence ellipsoid $\mathcal{E}_{t}\coloneqq\left\{\theta\in\mathbb{R}^{d}:\|\theta-\widehat{\theta}_{t}\|_{V_{t}}\leq\beta_{t}\right\}$ to represent $\mathcal{E}^{RLS}_{t,\delta^{\prime},\lambda_{\text{reg}}}$ and $\mathcal{E}^{PVT}_{t,\delta^{\prime},\lambda_{\text{reg}}}$, since the proof works for both of them.